One billion electron flows over a period of 1 microsec from A to B.determine the average current I1.

Grade:12th pass

2 Answers

Eshan
5 years ago
Dear student,

Current is the amount of charge flowing per unit time.

HenceI=\dfrac{ne}{t}=\dfrac{10^{9}\times 1.6\times 10^{-19}}{10^{-6}}A=0.16 mA
Gitanjali Rout
5 years ago
I=\dfrac{ne}{t}=\dfrac{10^{9}\times 1.6\times 10^{-19}}{10^{-6}}A=0.16 mACurrent is the amount of charge flowing per unit time.Hence
hence answer is 0.16
hope it is useful
thank you